40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Oct 27 15:57:27 2020, shruti, DailyProgress, Lab Work, psoma locking  
    Reply  Wed Oct 28 12:42:31 2020, shruti, DailyProgress, Lab Work, psoma locking  Setup2020Oct27.pdfpsoma_PLL.pdfpsoma_pll_freespace_intermediate.pdf
       Reply  Thu Nov 19 11:23:44 2020, shruti, Update, Lab Work, psoma locking  PLL_FS_sym.pdfMZ_PLL_FS_sym.pdfNewBeat.pdf
          Reply  Tue Nov 24 10:16:04 2020, shruti, Update, Lab Work, psoma locking  BeatSetup.pdfBeatOrNot.pdf
             Reply  Fri Nov 27 13:38:18 2020, Shruti, Update, Lab Work, psoma locking  Beat.mp4BeatZoomed.pdf
                Reply  Fri Nov 27 19:43:33 2020, Koji, Update, Lab Work, psoma locking  
Message ID: 2576     Entry time: Tue Oct 27 15:57:27 2020     Reply to this: 2577
Author: shruti 
Type: DailyProgress 
Category: Lab Work 
Subject: psoma locking  

Today I modified the optical setup with the aim of obtaining the beat between the two diode lasers for phase-locking.

I added pick-off polarizing beamsplitters with HWPs in each path for now (to be able to adjust their power) and mixed them at a 50/50 non-polarizing beam-splitter to eventually reach a Newfocus 1811 low noise PD.

I will add pictures and more details later.

 

ELOG V3.1.3-